OpenStudy (anonymous):

solve w^(4)-12w^(2)-2=0 what is the solution w=?

OpenStudy (anonymous):

X^2 - 12X - 2 = 0

OpenStudy (anonymous):

Can you do it?

OpenStudy (karatechopper):

course he can! after chlorophyll's step you need to ______. what goes in the blank?

OpenStudy (anonymous):

(12 +/- SQRT(144+8))/2 (12 +/- SQRT(152))/2 (12 + 12.33)/2 = 12.165 (12 - 12.33)/2 = -0.165

OpenStudy (anonymous):

That means: w^2 = 12.165 Since the other root is negative and a square of a number can't be negative. => w = SQRT(12.165) => w = 3.488

OpenStudy (anonymous):

solve the equation for u then replace w^2 with u

OpenStudy (anonymous):

Look at my solution. That is what I did.

OpenStudy (anonymous):



OpenStudy (anonymous):